# Nuisance parameters and elliptically symmetric distributions: a geometric approach to parametric and semiparametric efficiency

### v2

@@ III A short introduction to the RES distributions
In this section, we briefly recall the definition and the main properties of the RES distributions distributions. We will then present and discuss a “canonical parameterization” that can be built upon them. Many different, yet equivalent, representation of the RES family can be found in the literature [18, 19, 9, 10, 8]. Here, we adopt the approach and the notation introduced in [8]. Moreover, in the following, we will consider only the absolutely continuous case, i.e. we suppose that each distribution admits a density w.r.t. the Lebesgue measure on $\mathbb{R}^{m}$. Before moving on, we would like to underline that some of the outcomes discussed in the following sections have already been presented in [9, 10] by using a “semiparametric generalization” of the Le Cam theory on Local Asymptotically Normal (LAN) families of distributions [20], [21, Ch.6]. Our main goal here is to recast the problem of the statistical inference in RES distributions in the framework of the Hilbert spaces and present a systematic analysis based on purely geometrical concepts, such as those used in Lemmas II.1 and II.2.
@@ III-A Essentials on RES distributions
It is immediate to verify that the definition of elliptical density suffers from a lack of indentifiability for the couple $(\bm{\Sigma},g)$. Specifically, we can easily note that $RES_{m}(\bm{\mu},\bm{\Sigma},g(t))\equiv RES_{m}(\bm{\mu},c^{m/2}\bm{\Sigma},g(ct)),\forall c>0$. To avoid this ambiguity, we decided to put a constraint on the “functional form” of the density generator $g$. In particular, we force $g$ to belong to the following set:
@@
Note that, from the stochastic representation (31), the properties of $\mathbf{u}\sim\mathcal{U}(S_{\mathbb{R}}^{m-1})$ and the fact that $\mathcal{Q}$ is independent of $\mathbf{u}$, the constraint $E\{\mathcal{Q}\}=m$ implies that
@@
i.e., the scatter matrix can be directly interpreted as the usual covariance matrix.
@@
We note in passing that the most popular choices for the scale function $S(\cdot)$ are $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$, $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$ and $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$. It is worth emphasizing that, under A1, A2 and A3, and as a direct consequence of the implicit function theorem, the first top-left entry of $\mathbf{V}_{S}$, i.e. $\mathbf{V}_{S}]_{11}$, can (locally at $\mathbf{V}_{S}$) be expressed as function of the other entries. For example, if we choose $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$, then $\mathbf{V}_{S,0}]_{11}$ is trivially given by $\mathbf{V}_{S}]_{11}=1$. For $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$, we have that $\mathbf{V}_{S}]_{11}=m-\sum\nolimits_{i=2}^{m}[\mathbf{V}_{S}]_{ii}$. Lastly, for $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$ it can be easily shown that, using the Laplace’s expansion of the determinant along the first row of $\mathbf{V}_{S}$, $\mathbf{V}_{S}]_{11}$ can be recovered as $\mathbf{V}_{S}]_{11}=\frac{1}{C_{11}}\left(1-\sum\nolimits_{i=2}^{m}(-1)^{1+i}[\mathbf{V}_{S}]_{1i}C_{1i}\right)$ where $C_{ij}$ indicates the cofactor of $\mathbf{V}_{S}]_{ij}$.
As a consequence, the RES model can be parameterized in two different, yet equivalent, ways:
$\mathcal{P}_{\bm{\nu},g}=\left\{p_{X}(\mathbf{x}|\bm{\nu},g)=|\bm{\Sigma}|^{-1/2}g\left((\mathbf{x}-\bm{\mu})^{T}\bm{\Sigma}^{-1}(\mathbf{x}-\bm{\mu})\right);\bm{\nu}\in\Omega,g\in\overline{\mathcal{G}}\right\},$ (39)
@@
$\mathcal{P}_{\bm{\eta},g}=\left\{p_{X}(\mathbf{x}|\bm{\eta},g)=s^{-m/2}|\mathbf{V}_{S}|^{-1/2}g\left(s^{-1}(\mathbf{x}-\bm{\mu})^{T}\mathbf{V}_{S}^{-1}(\mathbf{x}-\bm{\mu})\right);\bm{\eta}\in\Phi,g\in\overline{\mathcal{G}}\right\}$ (40)
@@
$\bm{\nu}\triangleq(\bm{\mu}^{T},\mathrm{vecs}(\bm{\Sigma})^{T})^{T}\in\Omega\subseteq\mathbb{R}^{m}\times\mathrm{vecs}(\mathcal{S}_{m}^{\mathbb{R}}),$ (41)
@@
In both of the parameterizations, the infinite-dimensional parameter is the density generator $g\in\overline{\mathcal{G}}$.
@@
It is important to note that the parameterization in (40) in terms of $(\underline{\mathrm{vecs}}(\mathbf{V}_{S}),s)$ is of interest for both practical and theoretical reasons. From a practical point of view, in a large part of applications such as Principal Component Analysis (PCA), Canonical Correlation Analysis (CCA) and subspace-based methods, only a scaled version of the covariance matrix, i.e. $\mathbf{V}_{S}$, is of interest while the scale term $s$ can be considered as a nuisance parameter. On the other hand, from a theoretical viewpoint, the parameterization in (40) allows us to investigate the hidden and counter-intuitive relationships between the parametric and semiparametric efficiency on the shape matrix $\mathbf{V}_{S}$ and the scale $s$ as in the analysis reported in [9, 10]. For this reason, in the following we will refer to $\mathcal{P}_{\bm{\eta},g}$ as the canonical parameterization of the RES model.
III-B The model $\mathcal{P}_{\bm{\eta},g}$ and its parametric and semiparametric submodels
The main reason for adopting the $\mathcal{P}_{\bm{\eta},g}$ model as canonical is that it can be analyzed using the geometric tools introduced in the Sect. II. In particular, in analogy to the models introduced in (6), (7), (8) and (9), and by indicating as
$\bm{\gamma}=(\bm{\mu}^{T},\underline{\mathrm{vecs}}(\mathbf{V}_{S})^{T})^{T}\in\Gamma\subseteq\mathbb{R}^{m}\times\underline{\mathrm{vecs}}(\mathcal{S}_{m,S}^{\mathbb{R}})$ (43)
the finite dimensional parameter vector of interest, we can define the following parametric submodel of $\mathcal{P}_{\bm{\eta},g}$:
$\mathcal{P}_{\bm{\gamma}}=\left\{p_{X}(\mathbf{x}|\bm{\gamma},s_{0},g_{0}):\bm{\gamma}\in\Gamma\right\},$ (44)
$\mathcal{P}_{s}=\left\{p_{X}(\mathbf{x}|\bm{\gamma}_{0},s,g_{0}):s\in\mathbb{R}^{+}\right\},$ (45)
$\mathcal{P}_{\bm{\gamma},s}=\left\{p_{X}(\mathbf{x}|\bm{\gamma},s,g_{0}):\bm{\gamma}\in\Gamma,s\in\mathbb{R}^{+}\right\},$ (46)
along with the non-parametric model
$\mathcal{P}_{g}=\left\{p_{X}(\mathbf{x}|\bm{\gamma}_{0},s_{0},g):g\in\bar{\mathcal{G}}\right\}.$ (47)
where the scale $s\in\mathbb{R}^{+}$ and the density generator $g\in\bar{\mathcal{G}}$ play the role of finite and infinite-dimensional nuisance parameters respectively.
@@ III-B The model $\mathcal{P}_{\bm{\eta},g}$ and its parametric and semiparametric submodels
In the case of RES distributions, we can define an additional parametric model that is largely exploited in applications. In fact, since the density generator $g\in\bar{\mathcal{G}}$ is generally parameterized by a set of $p$ parameters, say $\bm{\zeta}\in\Delta\subseteq\mathbb{R}^{p}$ (for example the degrees of freedom of the $t$-distribution or the shape and scale parameters of the Generalized Gaussian distribution), one can suppose to know a priori the functional form of $g\in\bar{\mathcal{G}}$ up to its parameter vector $\bm{\zeta}\in\Delta$. This leads to the definition of the following parametric model:
@@
$\mathcal{P}_{\bm{\gamma},s,\bm{\zeta}}=\left\{p_{X}(\mathbf{x}|\bm{\gamma},s,\bm{\zeta}):\bm{\gamma}\in\Gamma,s\in\mathbb{R}^{+},\bm{\zeta}\in\Delta\right\}$ (48)
where
$p_{X}(\mathbf{x}|\bm{\gamma},s,\bm{\zeta})=s^{-m/2}|\mathbf{V}_{S}|^{-1/2}g_{\bm{\zeta}}\left(s^{-1}(\mathbf{x}-\bm{\mu})^{T}\mathbf{V}_{S}^{-1}(\mathbf{x}-\bm{\mu})\right)$ (49)
and where $\bm{\zeta}\in\Delta$ represents a second finite-dimensional nuisance term. This framework will allow us to apply Lemmas II.1 and II.2 to bring to light the counterintuitive links (in terms of efficiency) between the scale parameter $s$, the parameters $\bm{\zeta}$ of $g$ and $g$ itself. First, however, we must derive the three score vectors of the finite-dimensional parameters, that is $\mathbf{s}_{\bm{\gamma}_{0}}=(\mathbf{s}_{\bm{\mu}_{0}}^{T},\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}^{T})^{T}$, $s_{s_{0}}$ and $s_{\bm{\zeta}_{0}}$.
III-C The score vectors and the related FIM in $\mathcal{P}_{\bm{\eta},g}$
@@
Calculating the score vectors $\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}$ and $s_{s_{0}}$ directly in $\mathcal{P}_{\bm{\eta},g}$ is not an easy task due to the fact that the differentiable manifold $\mathcal{S}_{m,S}^{\mathbb{R}}$ in (38) is not a linear subspace of the set of real matrices $\mathbb{R}^{m\times m}$. On the contrary, since $\mathcal{S}_{m}^{\mathbb{R}}$ is indeed a linear subspace of $\mathbb{R}^{m\times m}$, the score vector $\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}$ for the scatter matrix $\bm{\Sigma}$ in the model $\mathcal{P}_{\bm{\nu},g}$ in (39) is easy to derive and well-known in the literature (see e.g. [22]). Fortunately, due to the differentability (under A1, A2 and A3) of $\mathcal{S}_{m,S}^{\mathbb{R}}$, we can define a diffeomorphism, say $\mathbf{w}$, between the canonical parameterisation of $\mathcal{P}_{\bm{\eta},g}$ and the one of $\mathcal{P}_{\bm{\nu},g}$. This allows us to easily pass from the score vectors of $\mathcal{P}_{\bm{\nu},g}$ to those of $\mathcal{P}_{\bm{\eta},g}$. Let us start by providing an explicit expression for such diffeomorphism.
@@
By recalling that $\bm{\Sigma}=s\mathbf{V}_{S}$, under A1, A2 and A3, we can move from the parameterization (40) to the one in (39) (and vice versa) by means of
@@
$\begin{split}\mathbf{w}:\;&\Phi\rightarrow\Omega\\ &\bm{\eta}=(\bm{\mu}^{T},\underline{\mathrm{vecs}}(\mathbf{V}_{S})^{T},s)^{T}\mapsto\mathbf{w}(\bm{\eta})=(\bm{\mu}^{T},s\cdot[\mathbf{V}_{S}]_{11},s\cdot\underline{\mathrm{vecs}}(\mathbf{V}_{S})^{T})^{T},\end{split}$ (50)
@@
where $s_{0}=S(\bm{\Sigma}_{0})$ and $\mathbf{V}_{S,0}=\bm{\Sigma}_{0}/s_{0}$. Furthermore, the 2nd diagonal component in (51), as shown in [10, Sect. 4] and, by using our notation in the Appendix .4, of this work is given by
@@
where $S([\mathbf{V}_{S}]_{11},\underline{\mathrm{vecs}}(\mathbf{V}_{S}))$ denotes the scale function applied to $\mathbf{V}_{S}$ reconstructed from $\mathbf{V}_{S}]_{11}$ and $\underline{\mathrm{vecs}}(\mathbf{V}_{S})$. For further reference, let us recast $\mathbf{J}[\mathbf{w}](\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0})$ as:
@@
The derivation of the gradient $\mathbf{J}[\mathbf{w}](\bm{\eta}_{0})$ in (51) can be also found in [10, 9] as function of the matrix $\mathbf{M}_{S}^{\mathbf{V}_{S}}$ that can be linked to our results through the relation:
@@
Due to the crucial importance of this matrix in the following sections, we collect in the Appendix .6 both its known properties (from [10, 9]) and some original ones, which will be extensively used in the rest of this paper.
Let us now introduce the score vector for $\bm{\nu}_{0}\in\Omega$ in the model $\mathcal{P}_{\bm{\nu},g}$ in (39) as $\mathbf{s}_{\bm{\nu}_{0}}=(\mathbf{s}_{\bm{\mu}_{0}}^{T},\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}^{T})^{T}$. By applying the rules of the differential matrix calculus detailed in [23, Ch. 8], the Kronecker product and the vec operator [23, Ch. 3], and the stochastic representation theorem (31), we get the following expressions of the scores [22, eqs. (34) and (37)]:
$\mathbf{s}_{\bm{\mu}_{0}}=_{d}\sqrt{\mathcal{Q}}\varphi_{0}(\mathcal{Q})\bm{\Sigma}^{-1/2}_{0}\mathbf{u},$ (56)
$\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}=_{d}2^{-1}\mathbf{D}_{m}^{T}\left[\bm{\Sigma}_{0}^{-1/2}\otimes\bm{\Sigma}_{0}^{-1/2}\right]\left[\mathcal{Q}\varphi_{0}(\mathcal{Q})\mathrm{vec}(\mathbf{u}\mathbf{u}^{T})-\mathrm{vec}(\mathbf{I}_{m})\right],$ (57)
where, following the notation adopted in [8], we defined:  [^2] To avoid confusion, it is important to note that the function $\varphi_{0}(t)$ is linked to the equivalent function $\psi_{0}(t)$ used in [22, 24] by the constant 2-2: $\varphi_{0}(t)=-2\psi_{0}(t)$.
$\varphi_{0}(t)\triangleq\frac{-2}{g_{0}(t)}\frac{dg_{0}(t)}{dt}.$ (58)
By using the diffeomorphism in (50) and its Jacobian in (51), the score vector for $\bm{\eta}_{0}\triangleq(\bm{\gamma}_{0}^{T},s_{0})^{T}\in\Phi$ in $\mathcal{P}_{\bm{\eta},g}$ can be obtained from the derived score vector $\mathbf{s}_{\bm{\nu}_{0}}$ in $\mathcal{P}_{\bm{\nu},g}$ as:
$\mathbf{s}_{\bm{\eta}_{0}}=[\mathbf{J}[\mathbf{w}](\bm{\eta}_{0})]^{T}\mathbf{s}_{\bm{\nu}_{0}}=\left(\begin{array}[]{c}\mathbf{s}_{\bm{\mu}_{0}}\\ \left[\mathbf{J}[\mathbf{w}](\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0})\right]^{T}\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}\end{array}\right).$ (59)
@@
$\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}=_{d}2^{-1}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}]\left(\mathcal{Q}\varphi_{0}(\mathcal{Q})\mathrm{vec}(\mathbf{u}\mathbf{u}^{T})-\mathrm{vec}(\mathbf{I}_{m})\right)$ (61)

### v3

@@ III A short introduction to the RES distributions
In this section, we briefly recall the definition and the main properties of the RES distributions distributions. Many different, yet equivalent, representation of the RES family can be found in the literature [20, 21, 10, 11, 9]. Here, we adopt the approach and the notation introduced in [9]. Moreover, in the following, we will consider only the absolutely continuous case, i.e. we suppose that each distribution admits a density w.r.t. the Lebesgue measure on $\mathbb{R}^{m}$. Before moving on, we would like to underline that some of the outcomes discussed in the following sections have already been presented in [10, 11] by using a “semiparametric generalization” of the Le Cam theory on Local Asymptotically Normal (LAN) families of distributions [22], [23, Ch.6]. Our main goal here is to recast the problem of the statistical inference in RES distributions in the framework of the Hilbert spaces and present a systematic analysis based on purely geometrical concepts, such as those used in Lemmas II.1 and II.2.
@@ III-A Essentials on RES distributions
It is immediate to verify that the definition of elliptical density suffers from a lack of indentifiability for the couple $(\bm{\Sigma},g)$. Specifically, we can easily note that $RES_{m}(\bm{\mu},\bm{\Sigma},g(t))\equiv RES_{m}(\bm{\mu},c^{m/2}\bm{\Sigma},g(ct)),\forall c>0$. To avoid this ambiguity, we may decide to put a constraint on the “functional form” of the density generator $g$. In particular, we force the density generator to belong to the following set:
@@
Note that, from the stochastic representation (31), the properties of $\mathbf{u}\sim\mathcal{U}(S_{\mathbb{R}}^{m-1})$ and the fact that $\mathcal{Q}$ is independent of $\mathbf{u}$, the constraint $E\{\mathcal{Q}\}=m$ is equivalent to:
@@
i.e., the scatter matrix can be directly interpreted as the usual covariance matrix. It is important to highlight that, as shown in [10, 11], it is possible to choose more general constraints that do not require the finiteness of any moment of the 2nd-order modular variate $\mathcal{Q}$, unlike the one in (34) used in this work.
@@
We note in passing that the most popular choices for the scale function $S(\cdot)$ are $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$, $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$ and $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$. It is worth emphasizing that, under A1, A2 and A3, and as a direct consequence of the implicit function theorem, the first top-left entry of $\mathbf{V}_{S}$, i.e. $\mathbf{V}_{S}]_{11}$, can (locally at $\mathbf{V}_{S}$) be expressed as function of the other entries. For example, if we choose $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$, then $\mathbf{V}_{S}]_{11}$ is trivially given by $\mathbf{V}_{S}]_{11}=1$. For $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$, we have that $\mathbf{V}_{S}]_{11}=m-\sum\nolimits_{i=2}^{m}[\mathbf{V}_{S}]_{ii}$. Lastly, for $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$ it can be easily shown that, using the Laplace’s expansion of the determinant along the first row of $\mathbf{V}_{S}$, $\mathbf{V}_{S}]_{11}$ can be recovered as $\mathbf{V}_{S}]_{11}=\frac{1}{C_{11}}\left(1-\sum\nolimits_{i=2}^{m}(-1)^{1+i}[\mathbf{V}_{S}]_{1i}C_{1i}\right)$ where $C_{ij}$ indicates the cofactor of $\mathbf{V}_{S}]_{ij}$.
As discussed in [10], the RES model can be parameterized in two different, yet equivalent, ways:
$\mathcal{P}_{\bm{\gamma},g}=\left\{p_{X}(\mathbf{x}|\bm{\gamma},g)=|\mathbf{V}_{S}|^{-1/2}g\left((\mathbf{x}-\bm{\mu})^{T}\mathbf{V}_{S}^{-1}(\mathbf{x}-\bm{\mu})\right);\bm{\gamma}\in\Gamma,g\in\mathcal{G}\right\},$ (39)
@@
$\mathcal{P}_{\bm{\eta},\bar{g}}=\left\{p_{X}(\mathbf{x}|\bm{\eta},\bar{g})=s^{-m/2}|\mathbf{V}_{S}|^{-1/2}\bar{g}\left(s^{-1}(\mathbf{x}-\bm{\mu})^{T}\mathbf{V}_{S}^{-1}(\mathbf{x}-\bm{\mu})\right);\bm{\eta}\in\Phi,\bar{g}\in\overline{\mathcal{G}}\right\}$ (40)
@@
$\bm{\gamma}\triangleq(\bm{\mu}^{T},\underline{\mathrm{vecs}}(\mathbf{V}_{S})^{T})^{T}\in\Gamma\subseteq\mathbb{R}^{m}\times\underline{\mathrm{vecs}}(\mathcal{S}_{m,S}^{\mathbb{R}}),$ (41)
@@
In both parameterizations, the infinite-dimensional parameter is the density generator, specifically its unconstrained version $g\in\mathcal{G}$ in (39) and the constrained one $\bar{g}\in\overline{\mathcal{G}}$ in (40). Moreover, for any scale function $S$, we have that and $g(t)$ and $\bar{g}(t)$ are related by $g(t)=s^{-m/2}\bar{g}(s^{-1}t)$, $t\in\mathbb{R}^{+}$.
@@
It is important to note that the two parameterizations in terms of $\underline{\mathrm{vecs}}(\mathbf{V}_{S})$ are of interest for both practical and theoretical reasons. From a practical standpoint, in many applications such as Principal Component Analysis (PCA), Canonical Correlation Analysis (CCA) and subspace-based methods, only a scaled version of the covariance matrix, i.e. $\mathbf{V}_{S}$, is of interest while the scale term $s$ can be treated as a nuisance parameter. Moreover, from a theoretical viewpoint, the parameterization in (40) allows us to investigate the hidden and counter-intuitive relationships between the parametric and semiparametric efficiency on the shape matrix $\mathbf{V}_{S}$ and the scale $s$.
III-B Parametric and non-parametric sub-models of $\mathcal{P}_{\bm{\gamma},g}$ and $\mathcal{P}_{\bm{\eta},\bar{g}}$
For further reference, we summarize below the parametric and non-parametric sub-models of the two RES-models in (39) and (40) that will be extensively used in the rest of this paper. For the sake of clarity, we will indicate with a superscript $u$ the “unconstrained” sub-models, i.e. the ones involving an unconstrained density generator $g\in\mathcal{G}$, and with a superscript $c$ the “constrained” sub-models, i.e. the ones involving a constrained $\bar{g}\in\overline{\mathcal{G}}$.
For the model $\mathcal{P}_{\bm{\eta},\bar{g}}$ in (40), we may identify three parametric sub-models as:
$\mathcal{P}_{\bm{\gamma}}^{c}=\left\{p_{X}(\mathbf{x}|\bm{\gamma},s_{0},\bar{g}_{0}):\bm{\gamma}\in\Gamma\right\},$ (43)
$\mathcal{P}^{c}_{s}=\left\{p_{X}(\mathbf{x}|\bm{\gamma}_{0},s,\bar{g}_{0}):s\in\mathbb{R}^{+}\right\},$ (44)
$\mathcal{P}^{c}_{\bm{\gamma},s}=\left\{p_{X}(\mathbf{x}|\bm{\gamma},s,\bar{g}_{0}):\bm{\gamma}\in\Gamma,s\in\mathbb{R}^{+}\right\},$ (45)
and a non-parametric sub-model as:
$\mathcal{P}^{c}_{\bar{g}}=\left\{p_{X}(\mathbf{x}|\bm{\gamma}_{0},s_{0},\bar{g}):\bar{g}\in\overline{\mathcal{G}}\right\}.$ (46)
@@
We may also define a parametric sub-model of $\mathcal{P}_{\bm{\eta},\bar{g}}$ that is largely exploited in applications. In fact, since the constrained density generator $\bar{g}\in\overline{\mathcal{G}}$ is generally parameterized by a set of $p$ parameters, say $\bm{\zeta}\in\Delta\subseteq\mathbb{R}^{p}$ (for example the degrees of freedom of the $t$-distribution or the shape parameter of the Generalized Gaussian distribution), one may assume that functional form of $\bar{g}\in\overline{\mathcal{G}}$ is known a-priori, up to its parameter vector $\bm{\zeta}\in\Delta$. This leads to the definition of the following parametric sub-model:
@@
$\mathcal{P}^{c}_{\bm{\gamma},s,\bm{\zeta}}=\left\{p_{X}(\mathbf{x}|\bm{\gamma},s,\bm{\zeta}):\bm{\gamma}\in\Gamma,s\in\mathbb{R}^{+},\bm{\zeta}\in\Delta\right\}$ (47)
where $p_{X}(\mathbf{x}|\bm{\gamma},s,\bm{\zeta})=s^{-m/2}|\mathbf{V}_{S}|^{-1/2}\bar{g}_{\bm{\zeta}}\left(s^{-1}(\mathbf{x}-\bm{\mu})^{T}\mathbf{V}_{S}^{-1}(\mathbf{x}-\bm{\mu})\right)$ where $\bar{g}_{\bm{\zeta}}\in\overline{\mathcal{G}}$ and $\bm{\zeta}\in\Delta$ represents a second finite-dimensional nuisance term.
For model $\mathcal{P}_{\bm{\gamma},g}$ in (39), we can identify the parametric sub-model
$\mathcal{P}^{u}_{\bm{\gamma}}=\left\{p_{X}(\mathbf{x}|\bm{\gamma},g_{0}):\bm{\gamma}\in\Gamma\right\}$ (48)
and an unconstrained non-parametric sub-model as:
$\mathcal{P}^{u}_{g}=\left\{p_{X}(\mathbf{x}|\bm{\gamma}_{0},g):g\in\mathcal{G}\right\}.$ (49)
This framework will allow us to apply Lemmas II.1 and II.2 to reveal the counter-intuitive links (in terms of efficiency) between the different RES sub-models. The first step, which enables us to analyze parametric and semiparametric efficiency in the RES model, is the derivation of the score vectors of the finite-dimensional parameters, i.e. the location vector $\bm{\mu}\in\mathbb{R}^{m}$, the shape matrix $\mathbf{V}_{S}\in\mathcal{S}_{m,S}^{\mathbb{R}}$ (or of its vectorized version) and the scale $s\in\mathbb{R}^{+}$.
III-C The score vectors $\mathbf{s}_{\bm{\mu}_{0}}$, $\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}$ and $s_{s_{0}}$ in $\mathcal{P}^{c}_{\bm{\gamma},s}$
@@
Calculating the score vectors $\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}$ and $s_{s_{0}}$ is not an easy task because the differentiable manifold $\mathcal{S}_{m,S}^{\mathbb{R}}$ in (38) is not a linear subspace of the set of real matrices $\mathbb{R}^{m\times m}$. On the contrary, since $\mathcal{S}_{m}^{\mathbb{R}}$ is indeed a linear subspace of $\mathbb{R}^{m\times m}$, the score vector $\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}$ for the scatter matrix $\bm{\Sigma}$ is easy to derive and well-known in the literature (see e.g. [24, 18]). Fortunately, due to the differentability (under A1, A2 and A3) of $\mathcal{S}_{m,S}^{\mathbb{R}}$, there exists a diffeomorphism, say $\mathbf{w}$, between $(\mathbf{V}_{S},s)$ and $\bm{\Sigma}$. Let us first provide an explicit expression for such diffeomorphism. It is important to highlight that most of the calculations that follow have already been derived in [10, 11]. In reporting them here, our aim is to provide a treatment consistent with the notations used in this article. We believe this will greatly facilitate reading and understanding.
@@
By recalling that $\bm{\Sigma}=s\mathbf{V}_{S}$, under A1, A2 and A3, we can move from the parameterization based on $(\bm{\mu},\mathbf{V}_{S},s)$ to the one based on $(\bm{\mu},\bm{\Sigma})$ and vice versa by means of
@@
$\begin{split}\mathbf{w}:\;&\Phi\rightarrow\Omega\subseteq\mathbb{R}^{m}\times\mathrm{vecs}(\mathcal{S}_{m}^{\mathbb{R}})\\ &\bm{\eta}=(\bm{\mu}^{T},\underline{\mathrm{vecs}}(\mathbf{V}_{S})^{T},s)^{T}\mapsto\mathbf{w}(\bm{\eta})=(\bm{\mu}^{T},s\cdot[\mathbf{V}_{S}]_{11},s\cdot\underline{\mathrm{vecs}}(\mathbf{V}_{S})^{T})^{T},\end{split}$ (50)
@@
where $s_{0}=S(\bm{\Sigma}_{0})$ and $\mathbf{V}_{S,0}=\bm{\Sigma}_{0}/s_{0}$. Furthermore, the 2nd diagonal component in (51), as shown in [11, Sect. 4] and, by using our notation in the Appendix .5, of this work is given by
@@
where $\bm{\nabla}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}$ indicates the “constrained” gradient w.r.t. $\underline{\mathrm{vecs}}(\mathbf{V}_{S})\in\underline{\mathrm{vecs}}(\mathcal{S}_{m,S}^{\mathbb{R}})$ and $S([\mathbf{V}_{S}]_{11},\underline{\mathrm{vecs}}(\mathbf{V}_{S}))$ denotes the scale function applied to $\mathbf{V}_{S}$ reconstructed from $\mathbf{V}_{S}]_{11}$ and $\underline{\mathrm{vecs}}(\mathbf{V}_{S})$. For further reference, let us recast $\mathbf{J}[\mathbf{w}](\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0})$ as:
@@
The derivation of the gradient $\mathbf{J}[\mathbf{w}](\bm{\eta}_{0})$ in (51) can be also found in [11, 10] as function of the matrix $\mathbf{M}_{S}^{\mathbf{V}_{S}}$ which can be linked to our results through the relation:
@@
Due to the crucial importance of this matrix in the subsequent sections, we collect in the Appendix .7 both its known properties (from [11, 10]) and some new ones, which will be extensively used in the reminder of the paper.
Let us now introduce the score vectors $\mathbf{s}_{\bm{\mu}_{0}}$ and $\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}$. By applying the rules of the differential matrix calculus detailed in [25, Ch. 8], the Kronecker product and the vec operator [25, Ch. 3], and the stochastic representation theorem (31), we get the following expressions of the scores ([24, eqs. (34) and (37)],[18, Sec. 5.3.2.1]):
$\mathbf{s}_{\bm{\mu}_{0}}=_{d}\sqrt{\mathcal{Q}}\bar{\varphi}_{0}(\mathcal{Q})\bm{\Sigma}^{-1/2}_{0}\mathbf{u},$ (56)
$\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}=_{d}2^{-1}\mathbf{D}_{m}^{T}[\bm{\Sigma}_{0}^{-1/2}\otimes\bm{\Sigma}_{0}^{-1/2}]\mathrm{vec}(\mathcal{Q}\bar{\varphi}_{0}(\mathcal{Q})\mathbf{u}\mathbf{u}^{T}-\mathbf{I}_{m}),$ (57)
where, following the notation adopted in [9], we defined:  [^2] To avoid confusion, it is important to note that the function $\bar{\varphi}_{0}(t)$ is linked to the equivalent function $\psi_{0}(t)$ used in [24, 26] by the constant 2-2: $\bar{\varphi}_{0}(t)=-2\psi_{0}(t)$.
$\bar{\varphi}_{0}(t)\triangleq\frac{-2}{\bar{g}_{0}(t)}\frac{d\bar{g}_{0}(t)}{dt}.$ (58)
By using the diffeomorphism in (50) and its Jacobian in (51), the score vector for $\bm{\eta}_{0}\triangleq(\bm{\gamma}_{0}^{T},s_{0})^{T}\in\Phi$ in $\mathcal{P}_{\bm{\eta},\bar{g}}$ in (40) can be obtained as:
$\mathbf{s}_{\bm{\eta}_{0}}=\left(\begin{array}[]{c}\mathbf{s}_{\bm{\mu}_{0}}\\ \left[\mathbf{J}[\mathbf{w}](\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0})\right]^{T}\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}\end{array}\right).$ (59)
@@
$\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}=_{d}2^{-1}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}]\mathrm{vec}(\mathcal{Q}\bar{\varphi}_{0}(\mathcal{Q})\mathbf{u}\mathbf{u}^{T}-\mathbf{I}_{m})$ (61)